# Amorphous Boron Nitride as an Ultrathin Copper Diffusion Barrier for Advanced Interconnects

### v2

@@ 4 Theoretical Investigation of -BN Diffusion Barriers
Note that, due to the effects of temperature on atomic motion, diffusivity is expected to increase with temperature (as seen in Fig.7-c), so that only a sharp increase of diffusivity (as seen in Fig. Fig.7-b and -c) can be safely interpreted as a sign of barrier failure. All samples, regardless the thickness and morphology, were able to prevent the Cu diffusion at low temperatures. 7-nm thick films prevent or largely mitigate Cu diffusion even at high temperatures (up to 1000 K). In contrast, 3-nm and 5-nm thick films were only successful when lower cooling rates were applied during their formation.
@@
As with all properties, the barrier performance of films largely depends on their morphology. Thus, we focused on the morphology of films to better understand the trends shown in Fig. 7. In the case of 7-nm thick films, even films formed with the highest high cooling rates exhibited relatively low levels of -hybridized atoms, high density, and fewer nanovoids near the boundaries compared to thinner films. While films cooled rapidly (CR25 K/ps) have some voids and -hybridized atoms leading to very limited Cu diffusion at near interface, slowly cooled films had a smooth surface without any voids close to the interface, therefore no Cu diffusion. The consistent increase in Cu diffusivity is largely due to increased atomic motion as the temperature increases.
@@
In contrast, the morphology of 3-nm and 5-nm thick films is significantly different than the 7-nm thick films. At higher cooling rates, the proportion of -hybridized bonds increased rapidly and voids were observed near the interface between -BN and Cu films. Moreover, the surface of rapidly cooled 3-nm and 5-nm thick samples were noticably rough. The potential energy per atom shows that these films are less stable than the 7-nm thick samples at the same cooling rates, as shown in Fig.12 in B, which can be attributed to larger number of -hybridized atoms and homonuclear bonds in thinner films. The lower density at high cooling rates (1.9-2.0 for 3-nm and 5-nm thick films, for 7-nm thick films at CR25 K/ps) along with the other morphological features mentioned here, indicates the existence of the voids in the films cooled rapidly. These voids and -hybridized atoms created pathways through the barrier which allowed Cu atoms to diffuse into the -BN. As Cu atoms close to the interface diffused into the -BN layer, B and N atoms also migrated into the Cu layer. This process accelerated the void formation and finally -BN barriers lost their structural integrity completely. Although providing visual representation of existing voids close to the interface is not easy, we present two 3-nm films which cooled at a rate of 100 K/ps (a) and 5 K/ps (b) in Fig. 8 after the annealing process. The black line shows the initial position of the interface. As shown in the Fig. 8, after the annealing process, the slow-cooled -BN barrier stopped Cu atoms from diffusing into the structure (only a few seemed to be diffused into the near interface area, which can be considered negligible). On the other hand, the fast-cooled -BN structure contained many small voids and had a worse stability. After the annealing at high temperatures, many Cu atoms diffused into the barrier, causing it to lose its structural integrity completely. Additionally, some B and N atoms also diffused into the Cu film, although in much smaller quantity compared to the diffused Cu atoms into the -BN film. While B and N atoms largely diffused into the regions close to the interface, Cu atoms diffused even deep into the -rm. We observed a similar process in other fast-cooled samples (CR: 25-100 K/ps) and 5-nm thick samples as well. However, it should be noted that 5-nm thick films managed to stop Cu diffusion at higher temperatures, however, fast-cooled 5-nm thick films were not able to prevent the Cu diffusion.
@@
These differences in morphology, caused by the different cooling protocols, demonstrate that barrier performance of the -BN films largely depends on the film quality. While films generated with cooling rates of 5 K/ps and 10 K/ps could stop the Cu diffusion until 1000 K, films with high coordination defects and voids onlymanaged to do the same until 500-600 K. These results suggest that ultrathin -BN thin films can serve as effective diffusion barriers and dielectric layer in integrated circuits using Cu wiring.
@@ 5 Experimental Investigation of -BN Diffusion Barriers
After we showed the potential of -BN films using theoretical models, we conducted an experimental investigation on the diffusion barrier performance of -BN films with the same thicknesses we performed our simulations. We fabricated Cu/aBN/Si structures and annealed them at 400 - 600 ℃ for 30 minutes. After annealing at 400 ℃, a Si substrate coated with a 3-nm thick aBN film did not form any copper silicide (Fig. 9-a), while a Si substrate not covered with aBN formed a substantial amount of copper silicide (Fig. 9-e). After annealing 3-nm thick aBN/Si at 500 ℃, very small copper silicide nucleations were observed on the Si surface (Fig. 9-b), and the formation of copper silicide crystals was confirmed after annealing at 600°C (Fig. 9-d). Fig. 9-c provides a brief summary of Cu diffusion barrier properties of 3-nm, 5-nm, and 7-nm thick aBN films after the 400 - 600 ℃ annealing process. The observation of crystallization by SEM is fine but not sensitive enough to determine Cu diffusion in the sub-silicide formation regime - that is - you could have Cu in the Si but SEM resolution is insufficient to detect or within the solubility limit of Cu in Si (Fig. 13 in C). The characteristics such as density or existence of defects of experimentally synthesized aBN films differ from the simulated aBN films and do not exhibit the same barrier performance as the simulations. However, there is a tendency for the barrier property to improve as the thickness of the aBN film increases.
@@ 6 Conclusions
We have investigated the diffusion barrier properties of ultrathin amorphous boron nitride (-BN) films against copper penetration by molecular dynamics and supported the simulations with experimental data. We focused on -BN films with thicknesses in the few nanometers range, and have analysed the morphologies, mechanical, diffusion barrier and optical properties of various -BN samples generated by molecular dynamics with machine learning trained interatomic potentials. The theoretical results agree with experimental barrier performance against Cu diffusion measured on PECVD-deposited 3-, 5- and 7-nm thick -BN films. Overall, the results confirm that amorphous boron-nitride compounds can be used as a dielectric barrier against Cu diffusion as a way to reduce the time delay in interconnects architectures.

### v3

@@ 4 Theoretical Investigation of -BN Diffusion Barriers
Note that, due to the effects of temperature on atomic motion, diffusivity is expected to increase with temperature (as seen in Fig.7-c), so that only a sharp increase of diffusivity (as seen in Fig.7-b and -c) can be safely interpreted as a sign of barrier failure. All samples, regardless the thickness and morphology, were able to prevent the Cu diffusion at low temperatures. 7-nm thick films prevent or largely mitigate Cu diffusion even at high temperatures (up to 1000 K). In contrast, 3-nm and 5-nm thick films were only successful when lower cooling rates were applied during their formation.
@@
As with all properties, the barrier performance of films largely depends on their morphology. Thus, we focused on the morphology of films to better understand the trends shown in Fig. 7. In the case of 7-nm thick films, even films formed with the highest high cooling rates exhibited relatively low levels of -hybridized atoms, high density, and fewer nanovoids near the boundaries compared to thinner films. While films cooled rapidly (CR25 K/ps) have some voids and -hybridized atoms leading to very limited Cu diffusion at near interface, slowly cooled films had a smooth surface without any voids close to the interface, therefore no Cu diffusion. The consistent increase in Cu diffusivity is largely due to increased atomic motion as the temperature increases. The superior diffusion barrier performance of 7-nm films can be attributed to their reduced defects and increased structural order, as described in Section 3.1, which allows them to better resist atomic migration even under less favourable conditions.
@@
In contrast, the morphology of 3-nm and 5-nm thick films is significantly different than the 7-nm thick films. At higher cooling rates, the proportion of -hybridized bonds increased rapidly and voids were observed near the interface between -BN and Cu films. Moreover, the surface of rapidly cooled 3-nm and 5-nm thick samples were noticeably rough. The potential energy per atom shows that these films are less stable than the 7-nm thick samples at the same cooling rates, as shown in Fig.12 in B, which can be attributed to larger number of -hybridized atoms and homonuclear bonds in thinner films. The lower density at high cooling rates (1.9-2.0 for 3-nm and 5-nm thick films, for 7-nm thick films at CR25 K/ps) along with the other morphological features mentioned here, indicates the existence of the voids in the films cooled rapidly. These voids and -hybridized atoms created pathways through the barrier which allowed Cu atoms to diffuse into the -BN. While 3-nm and 5-nm films are more susceptible to these defects, films generated at lower cooling rates demonstrated morphological stability, showing that quality, rather than thickness alone, determines performance. As Cu atoms close to the interface diffused into the -BN layer, B and N atoms also migrated into the Cu layer. This process accelerated the void formation and finally -BN barriers lost their structural integrity completely. Although providing visual representation of existing voids close to the interface is not easy, we present two 3-nm films which cooled at a rate of 100 K/ps (a) and 5 K/ps (b) in Fig. 8 after the annealing process. The black line shows the initial position of the interface. As shown in the Fig. 8, after the annealing process, the slow-cooled -BN barrier stopped Cu atoms from diffusing into the structure (only a few seemed to be diffused into the near interface area, which can be considered negligible). On the other hand, the fast-cooled -BN structure contained many small voids and had a worse stability. After the annealing at high temperatures, many Cu atoms diffused into the barrier, causing it to lose its structural integrity completely. Additionally, some B and N atoms also diffused into the Cu film, although in much smaller quantity compared to the diffused Cu atoms into the -BN film. While B and N atoms largely diffused into the regions close to the interface, Cu atoms diffused even deep into the -BN. We observed a similar process in other fast-cooled samples (CR: 25-100 K/ps) and 5-nm thick samples as well. However, it should be noted that 5-nm thick films managed to stop Cu diffusion at higher temperatures, however, fast-cooled 5-nm thick films were not able to prevent the Cu diffusion.
@@
These differences in morphology, caused by the different cooling protocols, demonstrate that the barrier performance of the -BN films largely depends on the film quality. While 3- and 5-nm thick films generated with cooling rates of 5 K/ps and 10 K/ps could stop the Cu diffusion until 1000 K, films with high coordination defects and voids (cooling rates greater than 50 K/ps) only managed to do the same until 500-600 K. These results suggest that ultrathin -BN films, when generated under optimised conditions, can perform comparably to thicker films as effective diffusion barriers and dielectric layers in integrated circuits using Cu wiring.
@@ 5 Experimental Investigation of -BN Diffusion Barriers
After we showed the potential of -BN films using theoretical models, we conducted an experimental investigation on the diffusion barrier performance of -BN films with the same thicknesses we performed our simulations. We fabricated Cu/-BN/Si structures and annealed them at 400 - 600 ℃ for 30 minutes. After annealing at 400 ℃, a Si substrate coated with a 3-nm thick -BN film did not form any copper silicide (Fig. 9-a), while a Si substrate not covered with -BN formed a substantial amount of copper silicide (Fig. 9-e). After annealing 3-nm thick -BN/Si at 500 ℃, very small copper silicide nucleations were observed on the Si surface (Fig. 9-b), and the formation of copper silicide crystals was confirmed after annealing at 600°C (Fig. 9-d). Fig. 9-c provides a brief summary of Cu diffusion barrier properties of 3-nm, 5-nm, and 7-nm thick -BN films after the 400 - 600 ℃ annealing process. The observation of crystallization by SEM is fine but not sensitive enough to determine Cu diffusion in the sub-silicide formation regime - that is - you could have Cu in the Si but SEM resolution is insufficient to detect or within the solubility limit of Cu in Si (Fig. 13 in C). The characteristics such as density or existence of defects of experimentally synthesized -BN films differ from the simulated -BN films and do not exhibit the same barrier performance as the simulations. However, there is a tendency for the barrier property to improve as the thickness of the -BN film increases.
@@ 6 Conclusions
We have investigated the diffusion barrier properties of ultrathin amorphous boron nitride (-BN) films against copper penetration by molecular dynamics and supported the simulations with experimental data. We focused on -BN films with thicknesses in the few nanometres range, and have analysed the morphologies, mechanical, diffusion barrier and dielectric properties of various -BN samples generated by molecular dynamics with machine learning trained interatomic potentials. The theoretical results agree with experimental barrier performance against Cu diffusion measured on PECVD-deposited 3-, 5- and 7-nm thick -BN films. Overall, the results confirm that amorphous boron-nitride compounds can be used as a dielectric barrier against Cu diffusion as a way to reduce the time delay in interconnects architectures. Further computational studies of -BN with larger thickness are however needed to explore (and find ways to mitigate) the effect of crystallisation observed in experiments, as it could could jeopardize reaching the upper performance and integrability of this innovative advanced material into mainstream interconnects technologies. Improved modelling of dielectric properties, beyond the qualitative tight-binding model used here, are likewise needed to better assess the properties of the films for such applications. In addition to a better electronic structure description, a more precise modelling of the dielectric function in the MHz to GHz range (including vibrational and intraband effects) would be of importance for comparison with experiments and to analyse the behaviour of -BN films at device operation frequencies.